The lengths of pregnancies in a small rural village are normally distributed with a mean of 264 days and a standard deviation of 12 days.

In what range would you expect to find the middle 50% of most pregnancies?
Between and .

If you were to draw samples of size 38 from this population, in what range would you expect to find the middle 50% of most averages for the lengths of pregnancies in the sample?
Between and .

Enter your answers as numbers. Your answers should be accurate to 1 decimal places.
Answer #1

Solution:-

Given that,

mean = \mu = 264

standard deviation = \sigma = 12

Middle 50%

\alpha = 1 - 50%  

\alpha = 1 - 0.50 = 0.50

\alpha/2 = 0.25

1 - \alpha /2 = 1 - 0.25 = 0.75

Z\alpha/2 = Z0.25 = -0.674

Z 1 \alpha /2 = Z 0.75 = 0.674

Using z-score formula,

x = z * \sigma + \mu

x = -0.674 * 12 +264

x = 255.912

Using z-score formula,

x = z * \sigma + \mu

x = 0.674 * 12 + 264

x = 272.088

Between 256.9 and 272.1

( b )

n = 38

\muT = 264

\sigmaT = \sigma / \sqrt n = 12 / \sqrt 38 = 1.9467

Middle 50%

\alpha = 1 - 50%  

\alpha = 1 - 0.50 = 0.50

\alpha/2 = 0.25

1 - \alpha /2 = 1 - 0.25 = 0.75

Z\alpha/2 = Z0.25 = -0.674

Z 1 \alpha /2 = Z 0.75 = 0.674

Using z-score formula  

T = z * \sigma T + \mu T  

T = -0.674 * 1.9467 + 264

T = 262.69

Using z-score formula  

T = z * \sigma T + \mu T  

T = 0.674 * 1.9467 + 264

T = 265.31

Between 262.7 and 265.3